Selected balance sheet amounts for Nordstrom Inc. for four recent years follow.

($ millions)

Current Assets

Long-term Assets

Total Assets

Current Liabilities

Long-term Liabilities

Total Liabilities

Equity

2013

5,228

3,346

2,541

3,953

2,080

2014

5,224

4,021

4,005

6,805

2015

3,014

7,698

2,911

6,827

871

2016

3,242

4,616

3,029

3,959

870

Compute the missing balance sheet amounts for each of the four years.
Answer #1

($ millions)

Current Assets

Long-term Assets

Total Assets

Current Liabilities

Long-term Liabilities

Total Liabilities

Equity

                    2,013

$    5,228.00

$ 3,346.00

$ 8,574.00

$ 2,541.00

$   3,953.00

$   6,494.00

$      2,080.00

                    2,014

$    5,224.00

$ 4,021.00

$ 9,245.00

$ 2,800.00

$   4,005.00

$   6,805.00

$      2,440.00

                    2,015

$    3,014.00

$ 4,684.00

$ 7,698.00

$ 2,911.00

$   3,916.00

$   6,827.00

$         871.00

                    2,016

$    3,242.00

$ 4,616.00

$ 7,858.00

$ 3,029.00

$   3,959.00

$   6,988.00

$         870.00

($ millions)

Current Assets

Long-term Assets

Total Assets

Current Liabilities

Long-term Liabilities

Total Liabilities

Equity

2013

5228

3346

=5228+3346

2541

3953

=2541+3953

2080

2014

5224

4021

=4021+5224

=6805-4005

4005

6805

=9245-6805

2015

3014

=7698-3014

7698

2911

=6827-2911

6827

871

2016

3242

4616

=3242+4616

3029

3959

=3029+3959

870